OpenStudy (anonymous):

What is the solution to the proportion? A. 4.5 B. 8 C. 9 D. 36

OpenStudy (anonymous):

oops!

OpenStudy (anonymous):

one sec xD

OpenStudy (anonymous):

6/x=9/12 is the preposition.

OpenStudy (anonymous):

sorry :c

OpenStudy (seanstewart16):

np

OpenStudy (anonymous):

brb gonna get a soda c:

OpenStudy (seanstewart16):

k

OpenStudy (anonymous):

Back c:

OpenStudy (seanstewart16):

ok, its 8.

OpenStudy (seanstewart16):

do u know why its eight?

OpenStudy (anonymous):

no :(

OpenStudy (seanstewart16):

ok, simplify 9/12

OpenStudy (anonymous):

3/4?

OpenStudy (seanstewart16):

YES

OpenStudy (anonymous):

oh ;D

OpenStudy (anonymous):

this soda is a year old <.<

OpenStudy (seanstewart16):

ok, know we need to make 6/x = 3/4

OpenStudy (anonymous):

should i drink it <.<

OpenStudy (seanstewart16):

thats old

OpenStudy (seanstewart16):

is it past expiration?

OpenStudy (anonymous):

nov/9/15 ;-;

OpenStudy (seanstewart16):

should still be good

OpenStudy (anonymous):

okay c: xD

OpenStudy (anonymous):

tasetes fine :D

OpenStudy (anonymous):

*tastes

OpenStudy (seanstewart16):

we need to make 6/x = 3/4

OpenStudy (anonymous):

three goes into six :D

OpenStudy (anonymous):

and two goes into four :D

OpenStudy (anonymous):

3/2

OpenStudy (seanstewart16):

no

OpenStudy (anonymous):

ah :c

OpenStudy (seanstewart16):

flip it, make 3/4 = 6/x. what must you do to 3 to make it 6?

OpenStudy (anonymous):

3+3?... or 3x2?

OpenStudy (seanstewart16):

were multiplying, so 2

OpenStudy (anonymous):

k

OpenStudy (seanstewart16):

now, because we multiplied 3 by 2, we also have to do it to the 4

OpenStudy (anonymous):

8

OpenStudy (seanstewart16):

YES!

OpenStudy (seanstewart16):

which gives you....

OpenStudy (anonymous):

6/8

OpenStudy (seanstewart16):

YES. now, to make sure we didint do it wrong, simplify 6/8

OpenStudy (anonymous):

8?

OpenStudy (seanstewart16):

simplify it like you did 9/12

OpenStudy (anonymous):

3/4

OpenStudy (seanstewart16):

YES

OpenStudy (anonymous):



OpenStudy (anonymous):

:D!!!!!!

OpenStudy (seanstewart16):

and because 9/12 = 3/4, and 6/8 = 3/4, you knowthat they are the same fraction. That is how you know it is eight

OpenStudy (seanstewart16):

understand now?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

i have more questions lined up c:

OpenStudy (anonymous):

fwee medals c:

OpenStudy (seanstewart16):

good job dude

OpenStudy (anonymous):

ty :D